OpenStudy (jjuden):

If m varies directly with n, and m = 6 when n = 9, what is the value of n when m = 8?

OpenStudy (jjuden):

a) n = 24 b) n = 12 c) n = 16/3 d) n = 27/4

OpenStudy (anonymous):

m =kn 6 =k(9) 6/9=k 8 =(6/9)n 8•(9/6) =n 12 =n

OpenStudy (anonymous):

yup its 12